OpenStudy (anonymous):

solve log25(x/2) = 1/2

OpenStudy (anonymous):

is it \[\log_{25}\frac{x}{2}=\frac{1}{2}\]????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

As we know, \[\log_{a}b=c => b=a^c\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so substitute values

OpenStudy (anonymous):

\[\large \frac{x}{2}=25^{\frac{1}{2}}\]

OpenStudy (anonymous):

....

OpenStudy (anonymous):

\[\frac{x}{2}=5\]

OpenStudy (anonymous):

x/2 = 5 x = 10